OpenStudy (anonymous):

Explain how you would find the exact value of √2 OVER √8.

OpenStudy (amistre64):

sqrts can "merge" \[\sqrt{a}*\sqrt{b}=\sqrt{ab}\]

OpenStudy (amistre64):

they can merge over a division as well

OpenStudy (unklerhaukus):

factorise eight, a pull any factors that are squared out of the sqrareroot

OpenStudy (loser66):

@ashbash0338 do as unkle said, quite easy to get the answer

OpenStudy (anonymous):

Factorize root 8 as (root 2 times root 4). root 4 becomes 2. and root 2 in denominator cancelles out root in numerator. so you are left with ....